OpenStudy (anonymous):

4.7x-4.2-3.3x=0.7(2x-6) solve

OpenStudy (anonymous):

i got x=0 is this correct and is this infinitly many

OpenStudy (mathstudent55):

Use the distributive property on the right side and then simplify both sides.

OpenStudy (anonymous):

If you got x=0 then there's only 1 solution, that is that x must equal 0. In this case however, you get: 4.7x-4.2-3.3x=0.7(2x-6) 1.4x - 4.2 = 1.4x -4.2 -4.2 = -4.2 0=0 This is always true, so yes there are infinitely many solutions!